# Mathematical Foundations and Operational Integration of ML for Uncertainty-Aware Decision-Making

### v1

@@ Abstract
Effective Uncertainty Quantification (UQ) represents a key aspect for reliable deployment of Large Language Models (LLMs) in automated decision-making and beyond. Yet, for LLM generation with multiple choice structure, the state-of-the-art in UQ is still dominated by the naive baseline given by the maximum softmax score. To address this shortcoming, we demonstrate that taking a principled approach via Bayesian statistics leads to improved performance despite leveraging the simplest possible model, namely linear regression. More precisely, we propose to train multiple Bayesian linear models, each predicting the output of a layer given the output of the previous one. Based on the obtained layer-level posterior distributions, we infer the global uncertainty level of the LLM by identifying a sparse combination of distributional features, leading to an efficient UQ scheme. Numerical experiments on various LLMs show consistent improvement over state-of-the-art baselines.
@@ 2.1 Background
In uncertainty quantification, it is common kendall2017uncertainties to distinguish between epistemic uncertainty, which arises from limited knowledge or imperfect models and can in principle be reduced with more data or better modeling, and aleatoric uncertainty, which captures the inherent randomness of the data-generating process and is irreducible. In the context of LLMs, a related dichotomy has emerged between factual uncertainty, which concerns uncertainty about the truthfulness of generated content relative to external world knowledge, and semantic uncertainty, which captures ambiguity or multiplicity in the possible meanings and continuations of a prompt. While the former aligns closely with epistemic uncertainty, the latter reflects linguistic indeterminacy inherent to natural language and is thus more akin to aleatoric variability. We focus in this work on factual UQ given its major role in automated decision making. Hence, we measure the LLM uncertainty in its response to Multiple Choice Questions, to disentangle semantic and factual uncertainties.

### v2

@@ Abstract
Effective Uncertainty Quantification (UQ) represents a key aspect for reliable deployment of Large Language Models (LLMs) in automated decision-making and beyond. Yet, for LLM generation with structured output, the state-of-the-art in UQ is still dominated by the naive baseline given by the maximum softmax score. To address this shortcoming, we demonstrate that taking a principled approach via Bayesian statistics leads to improved performance despite leveraging the simplest possible model, namely linear regression. More precisely, we propose to train multiple Bayesian linear models, each predicting the output of a layer given the output of the previous one. Based on the obtained layer-level posterior distributions, we infer the global uncertainty level of the LLM by identifying a sparse combination of distributional features, leading to an efficient UQ scheme. Numerical experiments on various LLMs show consistent improvement over state-of-the-art baselines.
@@ 2.1 Background
In uncertainty quantification, it is common kendall2017uncertainties to distinguish between epistemic uncertainty, which arises from limited knowledge or imperfect models and can in principle be reduced with more data or better modeling, and aleatoric uncertainty, which captures the inherent randomness of the data-generating process and is irreducible. In the context of LLMs, a related dichotomy has emerged between factual uncertainty, which concerns uncertainty about the truthfulness of generated content relative to external world knowledge, and semantic uncertainty, which captures ambiguity or multiplicity in the possible meanings and continuations of a prompt. While the former aligns closely with epistemic uncertainty, the latter reflects linguistic indeterminacy inherent to natural language and is thus more akin to aleatoric variability. We focus in this work on factual UQ given its major role in automated decision making. Hence, we measure the LLM uncertainty in its response to Multiple Choice Questions as a first step, to disentangle semantic and factual uncertainties.
@@ 4.2 Results
4.3 Extension to Open-Ended Questions
Building on the framework of kadavath2022language , we recast open-ended question uncertainty estimation into a structured setting by prompting the LLM to first generate an answer and then evaluate its own correctness. Based on the model’s internal activations during this process, we predict a confidence score for the produced answer. We compare our approach against state-of-the-art methods for open-ended uncertainty estimation, namely Semantic Entropy (SE) farquhar2024detecting and reflexive methods from kadavath2022language , hereafter referred to as P(True). To reduce the computational and memory costs of training, we implement our method as a sparse probe on the model activations, denoted Probe. Our results show a substantial improvement in the accuracy of confidence estimation, consistent with trends observed in the structured setting.
Table 2: AUROC for the different methods on the TriviaQA dataset. For each LLM, the best value is in bold and the second best is underlined. .
Llama-3.1-8B Qwen3-8B Ministral-8B SmolLM3-3B
P(True) 0.7710 0.7442 0.8154 0.8265
P(True)-10 0.7592 0.7194 0.7823 0.8237
SE 0.8288 0.7905 0.8352 0.8132
Probe (Ours) 0.8773 0.8958 0.8837 0.8708
@@ 5 Discussion & Conclusion
Broader Impact
This work advances safe and reliable AI by improving Uncertainty Quantification (UQ) in Large Language Models (LLMs)—a core component of trustworthy deployment alongside robustness dakhmouche2025noise ; rezkellah2025machine ; dakhmouche2025network ; dakhmoucherobust , privacy dong2022gaussian ; sajadmanesh2023privacy ; sajadmanesh2024progap and fairness castro2025representative ; zhang2024data . By offering an efficient alternative to computationally expensive ensemble or deep Bayesian methods, this work contributes to democratizing access to robust AI safety tools while fostering responsible human–AI interaction and reducing overconfidence in automated systems.